OpenStudy (anonymous):

determine the length of the leg of a 45-45-90 triangle with a hypotenuse length of 15 inches

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 x^2+x^2 = 15^2 2x^2 = 225 Keep going to solve for x